# On De Giorgi’s lemma for variational interpolants

### v2

@@ 4.3 De Giorgi’s estimate for general dissipation potentials
Finally, the Mosco convergence follows from the results in [Att84, Sec. 3.4].
We are now ready to perform the proof of Theorem, which will follow the usual approach of looking at the case $\eta>0$, deriving suitable a priori bounds that are independent of $\eta$, and then performing the limit $\eta\to 0^{+}$.
Proof of Theorem 4.12: We bring into play the Moreau-Yosida approximations $({\mathcal{R}}_{\eta})_{\eta>0}$. Since for every $\eta>0$ each ${\mathcal{R}}_{\eta}$ is radially differentiable, Theorem 4.1 provides the De Giorgi identity
$\forall\,\sigma>0\,:\quad{\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta})+\sigma{\mathcal{R}}_{\eta}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}^{\,\eta}{-}u^{o})\big{)}+\int_{0}^{\sigma}\mathcal{C}_{\mathcal{R}_{\eta}}(u^{o},\rho;\widetilde{u}_{\rho}^{\,\eta})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o})\,.$ (4.19)
for every measurable variational interpolant $(0,\infty)\ni\sigma\to\widetilde{u}_{\sigma}^{\,\eta}\in J_{\sigma}(u^{o})$.
Step $1$: A priori bounds. Clearly, we have that ${\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta})\leq{\mathcal{E}}(u^{o})$ for every $\sigma,\,\eta>0$. We will now use the place-holder ${v}_{\sigma}^{\eta}:=\frac{1}{\sigma}(u_{\sigma,\eta}{-}u^{o})$. Using that ${\mathcal{E}}$ is bounded below by some constant $E_{0}$, we deduce from (4.19) that
$\sigma{\mathcal{R}}_{\eta}({v}_{\sigma}^{\eta})+E_{0}\leq\sigma{\mathcal{R}}_{\eta}({v}_{\sigma}^{\eta})+{\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta})\stackrel{{\scriptstyle(1)}}{{\leq}}{\mathcal{E}}(u^{o})\quad\text{for all }\sigma,\,\eta>0\,,$ (4.20)
where (1) derives from the minimality of $\widetilde{u}_{\sigma}^{\,\eta}$. Thus, it follows from (4.17) applied with $S:={\mathcal{E}}(u^{o})$ that there exists a constant $C_{S}>0$ such that $\sigma\|{v}_{\sigma}^{\eta}\|\leq C_{S}$. Now, recalling that ${v}_{\sigma}^{\eta}=\frac{1}{\sigma}(\widetilde{u}_{\sigma}^{\,\eta}{-}u^{o})$ we infer that $\|\widetilde{u}_{\sigma}^{\,\eta}\|\leq\sigma\|{v}_{\sigma}^{\eta}\|\ +\|u^{o}\|\leq C_{S}+\|u^{o}\|$. All in all, we proved that
$\exists\,\overline{E}>0\ \ \forall\,\eta\in(0,1]\,\ \forall\,\sigma\in(0,\sigma_{*}]\,:\qquad{\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta})+\|\widetilde{u}_{\sigma}^{\,\eta}\|\leq\overline{E}.$ (4.21)
Step $2$: Compactness. Let us fix a null sequence $\eta_{n}\downarrow 0$. By (4.21), we have $\widetilde{u}_{\sigma}^{\,\eta_{n}}\in\overline{B_{\overline{E}}(0)}$ for every $\sigma\in(0,\sigma_{*}]$ and $n\in{\mathbb{N}}$. Therefore, the multi-valued mapping
$\mathscr{U}:(0,\sigma_{*}]\rightrightarrows X,\qquad\mathscr{U}(\sigma):=\mathop{\mathrm{Ls}^{\mathrm{w}}}_{n\to\infty}\{\widetilde{u}_{\sigma}^{\,\eta_{n}}\}$ (4.22)
is well defined with non-empty weakly closed values, and $\mathscr{U}(\sigma)\subset\overline{B_{\overline{E}}(0)}$ for all $\sigma\in(0,\sigma_{*}]$. Here, the Kuratowski upper limit of the sets $\{\widetilde{u}_{\sigma}^{\,\eta_{n}}\}$ w.r.t. the weak topology of $X$ is defined by
$\forall\,u\in\mathop{\mathrm{Ls}^{\mathrm{w}}}_{n\to\infty}\{\widetilde{u}_{\sigma}^{\,\eta_{n}}\}\text{ if }\ \exists\,(\eta_{n_{k}})_{k\in{\mathbb{N}}}\,:\quad\widetilde{u}_{\sigma}^{\,\eta_{n_{k}}}\rightharpoonup u\quad\text{in }X\,.$ (4.23)
The results of [AuF09, Chap. 8] (see also [Mai05, Thm. A.5.4]) imply that $\mathscr{U}$ is measurable.
Recall $\mathcal{C}_{\mathcal{R}_{\eta_{n}}}(u^{o},\sigma;u)$ from (4.13) and introduce the function $\widehat{\mathcal{C}}_{{\mathcal{R}}}:(0,\sigma_{*}]\to[0,\infty)$ defined by
${\widehat{\mathcal{C}}}_{{\mathcal{R}}}(\sigma):=\liminf_{n\to\infty}\mathcal{C}_{\mathcal{R}_{\eta_{n}}}(u^{o},\sigma;\widetilde{u}_{\sigma}^{\,\eta_{n}})\,.$
It follows from the closedness condition (4.12) that for each fixed $n\in{\mathbb{N}}$ the functionals $(\sigma,u)\mapsto\mathcal{C}_{\mathcal{R}_{\eta_{n}}}(u^{o},\sigma;u)$ are sequentially weakly lower semicontinuous on $(0,\sigma_{*}]{\times}X$. Therefore, the mappings $(0,\sigma_{*}]\ni\sigma\mapsto\mathcal{C}_{\mathcal{R}_{\eta_{n}}}(u^{o},\sigma;\widetilde{u}_{\sigma}^{\,\eta_{n}})$ are measurable, since they are given by the composition of the measurable curves $\sigma\mapsto(\sigma,\widetilde{u}_{\sigma}^{\,\eta_{n}})$ and of the lsc functionals $\mathcal{C}_{\mathcal{R}_{\eta_{n}}}(u^{o},\cdot;\cdot)$. Hence, their lower limit $\widehat{\mathcal{C}}_{{\mathcal{R}}}$ is measurable.
Additionally, we define the value function $\widehat{\mathscr{C}}_{{\mathcal{R}}}\colon(0,\sigma_{*}]\to[0,\infty)$ (observe the different script font $\mathscr{C}$) via
$\widehat{\mathscr{C}}_{{\mathcal{R}}}(\sigma):=\min\big{\{}\,\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;{u}_{\sigma})\>\big{|}\>{u}_{\sigma}\in\mathscr{U}(\sigma)\,\big{\}}\,.$
Indeed, the $\inf$ is attained for every $\sigma>0$ since $\mathscr{U}(\sigma)$ is a weakly closed and bounded subset of $X$, and then one can resort to the closedness property (4.12). We may again check the measurability of $\widehat{\mathscr{C}}_{{\mathcal{R}}}$ by representing it as the composition of measurable mappings: namely, of the measurable multifuction $\sigma\mapsto(\sigma,\mathscr{U}(\sigma))$ and of the mapping defined by
$(\sigma,U)\mapsto\mathcal{V}_{{\mathcal{R}}}(\sigma,U):=$ $\min\big{\{}\,\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;v)\>\big{|}\>v\in U\,\big{\}}$
for all $\quad\text{for all }\sigma\in(0,\sigma_{*}]\text{ and }U\text{ \emph{ weakly closed} subset of }\overline{B_{\overline{E}}(0)}\subset X\,.$
In turn, the latter value function is measurable as it is lower semicontinuous w.r.t. convergence in $(0,\infty)$ for $\sigma$, and Kuratowski convergence (induced by the weak topology of $X$) for $U$; we leave the details to the reader, as the argument is analogous to those in the previous and following lines.
We observe that
${\widehat{\mathcal{C}}}_{\mathcal{R}}(\sigma)\geq\widehat{\mathscr{C}}_{{\mathcal{R}}}(\sigma)\qquad\text{for all }\sigma\in(0,\sigma_{*}]\,.$ (4.24)
Indeed, fix $\sigma\in(0,\sigma_{*}]$, let $(\eta_{n})_{n}$ be a (not relabeled) subsequence such that
${\widehat{\mathcal{C}}}_{\mathcal{R}}(\sigma)=\lim_{n\to\infty}\mathcal{C}_{\mathcal{R}_{\eta_{n}}}(u^{o},\sigma;\widetilde{u}_{\sigma}^{\,\eta_{n}})=\lim_{n\to\infty}{\mathcal{R}}_{\eta_{n}}^{*}({-}\xi_{n})$
where $\xi_{n}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma}^{\,\eta_{n}})$ attains the optimal value of ${\mathcal{R}}_{\eta_{n}}^{*}({-}\xi)$, cf. Lemma 4.11. Thus, $\sup_{n}{\mathcal{R}}_{\eta_{n}}^{*}({-}\xi_{n})<\infty$, so that by the coercivity of ${\mathcal{R}}^{*}$ the sequence $(\xi_{n})_{n}$ is bounded in $X^{*}$ and hence it admits a not relabeled subsequence weakly converging as $n\to\infty$ to some $\overline{\xi}$. Since the dual dissipation potentials ${\mathcal{R}}_{\eta_{n}}^{*}$ Mosco-converge to ${\mathcal{R}}^{*}$, we have that $\liminf_{n\to\infty}{\mathcal{R}}_{\eta_{n}}^{*}({-}\xi_{n})\geq{\mathcal{R}}^{*}({-}\overline{\xi})\,.$ Now, $(\widetilde{u}_{\sigma}^{\,\eta_{n}})_{n}$ is bounded in $X$, hence there exists $\overline{u}$ such that, along a further subsequence, $\widetilde{u}_{\sigma}^{\,\eta_{n}}\rightharpoonup\overline{u}$. Therefore, $\overline{u}\in\mathscr{U}(\sigma)$ and, by the closedness property (4.12), we find that $\overline{\xi}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\overline{u})$. All in all, we have found that
$\lim_{n\to\infty}{\mathcal{R}}_{\eta_{n}}^{*}({-}\xi_{n})$ $\geq{\mathcal{R}}^{*}({-}\overline{\xi})\geq\min\big{\{}\,{\mathcal{R}}^{*}({-}\xi)\>\big{|}\>\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u),\ u\in\mathscr{U}(\sigma)\,\big{\}}\,,$
and (4.24) follows.
Step $3$: Passing to the limit in (4.19). Let us fix $\sigma\in(0,\sigma_{*}]$ and an arbitrary element $u\in\mathscr{U}(\sigma)$. By (4.23), there exists a sequence $(\eta_{n_{k}})_{k}$ (whose dependence on $\sigma$ and $u$ is not highlighted), such that (4.23) holds. Therefore, by the weak lower semicontinuity of ${\mathcal{E}}$ and by the Mosco-convergence of ${\mathcal{R}}_{\eta_{n_{k}}}$ to ${\mathcal{R}}$, we have
${\mathcal{E}}(u)\leq\liminf_{k\to\infty}{\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta_{n_{k}}})\,,$ (4.25)
$\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(u{-}u^{o})\big{)}\leq\liminf_{k\to\infty}\sigma{\mathcal{R}}_{\eta_{n_{k}}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}^{\,\eta_{n_{k}}}{-}u^{o})\big{)}\,.$
Finally, we have
@@
$\leq\liminf_{k\to\infty}\int_{0}^{\sigma}{\mathcal{R}}_{\eta_{n_{k}}}^{*}({-}\widetilde{\xi}_{\rho}^{\,\eta_{n_{k}}})\;\!\mathrm{d}\rho\,,$
where the last estimate follows from Fatou’s Lemma. Since the lower semicontinuity estimates (4.25) hold for every $u\in\mathscr{U}(\sigma)$, we arrive at the following estimate:
@@
$\forall\,\sigma\in(0,\sigma_{*}]\,:\quad\sup_{u\in\mathscr{U}(\sigma)}\left({\mathcal{E}}(u){+}\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(u{-}u^{o})\big{)}\right)+\int_{0}^{\sigma}\widehat{\mathscr{C}}_{{\mathcal{R}}}(\rho)\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{o})\,.$ (4.26)
Step $4$: Measurable selections and conclusion of the proof. We now resort to the multivalued mapping $\mathfrak{U}_{{\mathcal{R}}}:(0,\sigma_{*}]{\times}X\rightrightarrows X{\times}X^{*}$ defined by
$\mathfrak{U}_{{\mathcal{R}}}(\sigma,u):=\{u\}\times\mathfrak{A}_{{\mathcal{R}}}(\sigma,u)=\{u\}\times\mathrm{argmin}\big{\{}\,{\mathcal{R}}^{*}({-}\xi)\>\big{|}\>\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u)\,\big{\}}$
(cf. (4.15)). Clearly, $\mathfrak{U}_{{\mathcal{R}}}$ is also upper semicontinuous w.r.t. convergence in ${\mathbb{R}}$ and weak convergence in $X$, thus the composed multivalued mapping
$(0,\sigma_{*}]\ni\sigma\mapsto\mathfrak{U}_{{\mathcal{R}}}(\sigma,\mathscr{U}(\sigma))\subset X{\times}X^{*}$
is measurable. Then, applying [CaV77, Thm. 3.22] we conclude that for every $\sigma\in(0,\sigma_{*}]$ there exists a measurable selection

### v3

@@ 4.3 De Giorgi estimates for general dissipation potentials
The Mosco convergence follows from the results in [Att84, Sec. 3.4]. Finally, in order to show (4.23), we note that $\xi_{\eta}\in\partial{\mathcal{R}}_{\eta}(v_{\eta})$ equivalent to
$\big\langle\xi_{\eta},v_{\eta}\big\rangle={\mathcal{R}}_{\eta}\big(v_{\eta}\big)+{\mathcal{R}}^{*}_{\eta}(\xi_{\eta})={\mathcal{R}}_{\eta_{n}}\big(v_{\eta}\big)+{\mathcal{R}}^{*}(\xi_{\eta})+\tfrac{\eta}{2}\|\xi_{\eta}\|_{X^{*}}^{2}.$
Using that $v_{\eta}\to v$ and $\xi_{\eta}\rightharpoonup\xi$ we can pass to the limit on the left-hand side. On the right-hand side we can pass to the liminf and obtain $\langle\xi,v\rangle\geq{\mathcal{R}}(v)+{\mathcal{R}}^{*}(\xi)$, which yields the desired result $\xi\in\partial{\mathcal{R}}(v)$.
The following result shows that Hypothesis 4.11 can be naturally deduced if the conditioned subdifferential $\partial_{{\mathcal{R}}}{\mathcal{E}}$ has a suitable closedness property. A sufficient condition for such closedness involves the following property, which we define in general: We say that a functional ${\mathcal{F}}:X\to(-\infty,\infty]$ has the weak-implies-strong property ($\mathrm{WIS}$, for short), if the following implication holds:
$u_{k}\overset{X}{\rightharpoonup}u\text{ (weakly) }\text{ and }{\mathcal{F}}(u_{k})\to{\mathcal{F}}(u)<\infty\quad\Longrightarrow\quad u_{k}\overset{X}{\to}u\text{ (strongly)}.$ (4.25)
This property is certainly true if $\mathcal{F}$} has compact sublevels, however it is also true for uniformly convex functions, or for functionals ${\mathcal{F}}(u)=g(\|u\|)$ if $g$ is strictly increasing and $\|\cdot\|$ is a uniformly convex norm. Thus, the conditions in Prop. 4.14 are sufficiently easy to check, and indicate the potential for generalizations.
Proposition 4.14 (Sufficient conditions for (4.19))
Let the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypothesis 3.1 and fix $\sigma_{*}>0$ and $u^{\circ}\in\mathop{\mathrm{dom}}({\mathcal{E}})$.
(A) If for all $\sigma\in(0,\sigma_{*})$, the operator $\partial_{{\mathcal{R}}}{\mathcal{E}}:\mathrm{Graph}(\bm{J})\rightrightarrows X^{*}$ satisfies
@@
then the lower semi-continuity property (4.19) is satisfied.
(B) If $\mathcal{E}$} or $\mathcal{R}$} have the $\mathrm{WIS}$ property and $\mathrm{dom}({\mathcal{R}})=X$ with $\mathcal{R}$$^{*}$ having superlinear growth at infinity, then condition (4.29) holds.
Proof. For (A) the proof is straightforward. We note that for a sequence $(\eta_{n},u_{n})\rightharpoonup(0,u)$ with $\alpha:=\liminf_{n\to\infty}\mathcal{C}_{{\mathcal{R}}_{\eta_{n}}}(\sigma;u_{\eta})<\infty$ we find $\xi_{n}\in\partial_{{\mathcal{R}}_{\eta_{n}}}{\mathcal{E}}(\sigma;u_{n})\subset X^{*}$ with ${\mathcal{R}}_{\eta_{n}}^{*}(\xi_{n})\leq\mathcal{C}_{{\mathcal{R}}_{\eta_{n}}}(\sigma;u_{\eta})+\tfrac{\displaystyle 1}{\displaystyle n}$. In turn, since ${\mathcal{R}}^{*}(\xi_{n})\leq{\mathcal{R}}_{\eta_{n}}^{*}(\xi_{n})$, by the coercivity property (3.5) we conclude that $\|\xi_{n}\|_{X^{*}}\leq C<\infty$.
Thus, after extracting a suitable subsequence (not relabeled) we may assume $\xi_{n}\rightharpoonup\xi$ and the closedness property (4.29) implies $\xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;u)$. Now, using the weak lower semicontinuity of $\mathcal{R}$$^{*}$ we have
$\mathcal{C}_{{\mathcal{R}}}(\sigma;u)$ $\leq{\mathcal{R}}^{*}(\xi)\leq\liminf_{n\to\infty}{\mathcal{R}}^{*}_{\eta_{n}}(\xi_{n})=\liminf_{n\to\infty}\big(\mathcal{C}_{{\mathcal{R}}_{\eta_{n}}}(\sigma;u_{\eta})+\tfrac{\displaystyle 1}{\displaystyle n}\big)=\alpha,$
which is the desired result of part (A).
For part (B), let $(\eta_{n},u_{n},\xi_{n})$ be as in (4.29). We use that $u_{n}\in\bm{J}_{\eta_{n}}(\sigma)$ and $u_{n}\rightharpoonup u$ already imply
$\Phi_{\sigma}^{\eta_{n}}(u_{n})=\sigma{\mathcal{R}}_{\eta_{n}}\big(\tfrac{\displaystyle 1}{\displaystyle\sigma}(u_{n}{-}u^{\circ})\big)+{\mathcal{E}}(u_{\eta_{n}})\longrightarrow\Phi_{\sigma}(u)\,.$ (4.30)
Indeed, the $\liminf$-estimate follows by (4.18), whereas for the $\limsup$ we observe that $\Phi_{\sigma}^{\eta_{n}}(u_{n})\leq\Phi_{\sigma}^{\eta_{n}}(u)\leq\Phi_{\sigma}(u)$, where the first inequality follows from $u_{\eta_{n}}\in\bm{J}_{\eta_{n}}(\sigma)$ and the second one from ${\mathcal{R}}_{\eta_{n}}\leq{\mathcal{R}}$. We also have the individual $\liminf$-estimates
@@
${\mathcal{R}}(\frac{1}{\sigma}(u{-}u^{\circ}))\leq\liminf_{n\to\infty}{\mathcal{R}}_{\eta_{n}}(\frac{1}{\sigma}(u_{n}{-}u^{\circ}))\quad\text{and}\quad{\mathcal{E}}(u)\leq\liminf_{n\to\infty}{\mathcal{E}}(u_{n}),$ (4.31)
the first one by the Mosco-convergence of ${\mathcal{R}}_{\eta_{n}}$ to $\mathcal{R}$}, and the second one by the weak lower semicontinuity of $\mathcal{E}$}. Combining (4.30) and (4.31) we see that both liminfs are limits and that the inequalities must be equalities. Hence, by assumed $\mathrm{WIS}$ property for either $\mathcal{R}$}, or $\mathcal{E}$}, we find the strong convergence $u_{n}\to u$. Now, we need to take pass to the limit, as $n\to\infty$, in both relations: $\xi_{n}\in\partial{\mathcal{E}}(u_{n})$ and $\xi_{n}\in\partial{\mathcal{R}}_{\eta_{n}}\big(\frac{1}{\sigma}(u_{n}{-}u^{\circ})\big)$. For the first relation, we resort to the closedness property (3.4d) for $\partial{\mathcal{E}}$ (observe that $u_{n}$ lies in an energy sublevel since ${\mathcal{E}}(u_{n})\leq\Phi_{\sigma}^{\eta_{n}}(u_{n})\leq{\mathcal{E}}(u^{\circ})$). For the second relation, we exploit the gained strong convergence $\frac{1}{\sigma}(u_{n}{-}u^{\circ})\to\frac{1}{\sigma}(u{-}u^{\circ})$ and property (4.23). All in all, we conclude $\xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;u)$, and (4.29) is established. This proves part (B).
4.4 Proof of Theorem 4.12
The first tool in the proof of Theorem 4.12 are the previously recapped properties of the Moreau-Yosida approximations ${\mathcal{R}}_{\eta}$. The second tool is the following standard result on the existence of measurable selections for Kuratowski upper limits which, like Lemma 4.3, also draws on [CaV77].
Proposition 4.15 (Measurable selection)
Consider a sequence of $(u^{k})_{k\in{\mathbb{N}}}$ of measurable functions $(0,\sigma_{*})\ni\sigma\mapsto u^{k}_{\sigma}\in M$, where $M$ is a compact metric space.
Then, there exists a measurable selection $\sigma\mapsto\widetilde{u}_{\sigma}$ of the set of all limit points of $(u^{k}_{\sigma})_{k}$, namely for all $\sigma\in(0,\sigma_{*})$ there exists a subsequence $(k^{\sigma}_{n})_{n\in{\mathbb{N}}}$ with $k^{\sigma}_{n}\to\infty$ and $u^{k^{\sigma}_{n}}_{\sigma}\to\widetilde{u}_{\sigma}$ in $M$.
Furthermore, if ${\mathcal{A}}:(0,\sigma_{*}){\times}M\to[0,\infty]$ is a measurable function such that ${\mathcal{A}}(\sigma,\cdot)$ is lower semicontinuous for all $\sigma$, then the measurable selection $\widetilde{u}$ can be chosen such that additionally
${\mathcal{A}}(\sigma,\widetilde{u}_{\sigma})\leq A(\sigma):=\liminf_{k\to\infty}{\mathcal{A}}(\sigma,u^{k}_{\sigma})\quad\text{for all }(0,\sigma_{*}).$
Proof. As $M$ is compact, we can define the multivalued mapping
$\mathscr{U}:(0,\sigma_{*})\rightrightarrows M,\qquad\mathscr{U}(\sigma):=\mathop{\mathrm{Ls}}_{k\to\infty}\{u^{k}_{\sigma}\},$ (4.32)
where $\mathop{\mathrm{Ls}}$ is the set of all limit points (in the sense of a Kuratowski upper limit, cf. (2.10b))
$w\in\mathop{\mathrm{Ls}}_{k\to\infty}\{u^{k}_{\sigma}\}\quad\Longleftrightarrow\quad\exists\,(k_{n})_{n\in{\mathbb{N}}}:\ k_{n}\to\infty\text{ and }u^{k_{n}}_{\sigma}\to w.$ (4.33)
Hence, each $\mathscr{U}(\sigma)$ is well defined, non-empty and closed. The results of [AuF90, Thm. 8.2.5] imply that $\mathscr{U}$ is a measurable multivalued mapping. Now, [CaV77, Thm. 3.22] grants the existence of a measurable selection for $\mathscr{U}$, i.e. a function $\sigma\mapsto\widetilde{u}_{\sigma}$ with $\widetilde{u}_{\sigma}\in\mathscr{U}(\sigma)$ for all $\sigma\in(0,\sigma_{*})$.
For the second part of the statement involving $\mathcal{A}$} we define
$\mathscr{U}_{\mathcal{A}}(\sigma):=\{\,u\in\mathscr{U}(\sigma)\>|\>{\mathcal{A}}(\sigma,u)\leq A(\sigma)\,\}.$
By the definition of $\mathscr{U}_{\mathcal{A}}$ and the lsc of $\mathcal{A}$} we know have $\mathscr{U}_{\mathcal{A}}(\sigma)\neq\emptyset$ for all $\sigma$. Hence, $\mathscr{U}_{\mathcal{A}}:(0,\sigma_{*})\rightrightarrows M$ is a measurable multivalued mapping with non-empty and closed values, and we can apply the selection theorem as above, see also [AuF90, Thm. 8.1.4].
We are now ready to carry out the proof of Theorem 4.12 which provides the De Giorgi simple and improved estimate for general dissipation potentials $\mathcal{R}$}. For this, consider variational integrands for the regularized $\mathrm{gBGS}$ with dissipation potential ${\mathcal{R}}_{\eta}$.
Proof of Theorem 4.12.
Step 1: Regularization. For $\eta\in(0,1]$ we consider the dissipation potentials ${\mathcal{R}}_{\eta}$, which are radially differentiable. Hence, by Theorem 4.2 there exist variational interpolants $\sigma\mapsto\widetilde{u}^{\eta}_{\sigma}$ such that the De Giorgi identity holds
@@
$\forall\,\sigma>0\,:\quad{\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta})+\sigma{\mathcal{R}}_{\eta}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}^{\,\eta}{-}u^{\circ})\big)+\int_{0}^{\sigma}\mathcal{C}_{{\mathcal{R}}_{\eta}}(\rho;\widetilde{u}_{\rho}^{\,\eta})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ})\,.$ (4.34)
In particular, we have $\widetilde{u}_{\sigma}^{\,\eta}\in\bm{J}_{\eta}(\sigma)$, and $\mathcal{C}_{{\mathcal{R}}_{\eta_{n}}}(\cdot;\cdot)$ is defined in (3.13).
Step 2: A priori bounds. Clearly, we have that ${\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta})\leq{\mathcal{E}}(u^{\circ})$ for all $\sigma$ and $\eta$. We will now use the place-holder ${v}_{\sigma}^{\eta}:=\frac{1}{\sigma}(u_{\sigma,\eta}{-}u^{\circ})$. Recalling that $\mathcal{E}$} is bounded below by some constant $E_{0}$, we deduce from (4.34) that